# A Secure Estimator with Gaussian Bernoulli Mixture Model

### v1

@@ IV-B Estimation update for the disordered new observation
A basic approach is to use Lemma 1 to calculate the estimation of the state sequence under $y_{\mathcal{O}^{*}}$. In detail, we can utilize the known Kalman gain $\{K_{t}|t<k\}$ under $y_{\mathcal{O}}$ and recompute the rest process under $y_{\mathcal{O}^{*}}$, as illustrated in Fig. 5.
However, employing a direct method to compute the optimal estimate reveals imitations in computational cost. When $|y_{\mathcal{O}}|$ is large, especially in systems with multiple sensors, the direct method requires computing the matrix inverse $O(|y_{\mathcal{O}}|)$ times, which consumes significant computational resources. Furthermore, as the observation set expands, the optimal state estimate approaches the real state more closely. Nevertheless, adding new observations to a larger $y_{\mathcal{O}}$ requires an increasing number of calculation steps to update the estimate, especially in multi-sensor systems. This indicates that using direct method to update cannot fully leverage the information from the previous estimation, and thus causes a waste of computational resources.
Figure 5: Recalculation steps for estimation update. When a new observation $y_{k}$ is inserted, the whole Kalman gain after time $k$ and the smoother gain for the entire sequence needs to be recalculated.
V Iterative approach for state estimation
In section IV, we discussed how to find the closed-form solution of equation (8), but its computational efficiency is poor. To speed up the solution and update process, we propose an iterative method for state sequence estimation that allows for rapid convergence to the optimal estimate.
The core idea of the iterative method is to fully utilize the previous estimates as the initial state in the iteration. By constructing a composite convex optimization problem, we design a proximal gradient descent algorithm to avoid redundant computations. Furthermore, we derive the convergence rate and the relationship between the iteration count and the reliability of estimates.
V-A The estimation of state sequence with partial observations
With the observation set $y_{\mathcal{O}}$, our objective is to optimize the function $\mathcal{L}_{\mathcal{O}}(\hat{x}_{0:N})$ in equation (8). The objective function incorporates two goals: first, the state estimates $\hat{x}_{0:N}$ should match the observations $\hat{y}_{\mathcal{O}}$; second, the state estimates $\hat{x}_{0:N}$ should conform to the known structure of the system. Therefore, $\mathcal{L}_{\mathcal{O}}$ can be considered as the sum of two functions:
@@ V-A The estimation of state sequence with partial observations
We can represent each $\hat{x}_{i}$ using a column vector $\hat{X}=col\{\hat{x}_{0|\mathcal{O}},\cdots,\hat{x}_{N|\mathcal{O}}\}$. Next, we can construct a sparse matrix $\tilde{\mathcal{A}}$, which captures the relationship between the states:
$\begin{bmatrix}\hat{x}_{0}\\ \hat{x}_{1}-A\hat{x}_{0}\\ \hat{x}_{2}-A\hat{x}_{1}\\ \vdots\\ \hat{x}_{N}-A\hat{x}_{N-1}\end{bmatrix}=\begin{bmatrix}I&0&0&\cdots&0\\ -A&I&0&\cdots&0\\ 0&-A&I&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&I\end{bmatrix}\begin{bmatrix}\hat{x}_{0}\\ \hat{x}_{1}\\ \hat{x}_{2}\\ \vdots\\ \hat{x}_{N}\end{bmatrix}=\tilde{\mathcal{A}}\hat{X}.$ (12)
@@
$\tilde{P}=\begin{bmatrix}P_{0}&0&0&\cdots&0\\ 0&Q&0&\cdots&0\\ 0&0&Q&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&Q\end{bmatrix}.$
With the help of equation (13), we convert Problem 2 to a composite convex optimization problem.
@@
$\begin{array}[]{clll}\displaystyle\arg\min_{\hat{X}}&~{}\mathcal{L}_{\mathcal{O}}=f_{\mathcal{O}}(\hat{X})+g(\hat{X})\end{array}$ (15)
@@
$f_{\mathcal{O}}(\hat{X})$ $=\sum_{i\in\mathcal{O}}\|y_{i}-\tilde{C}_{i}\hat{X}\|^{2}_{R^{-1}},$
$g(\hat{X})$ $=\|\hat{X}\|_{H}^{2},\quad H={\tilde{\mathcal{A}}^{T}\tilde{P}^{-1}\tilde{\mathcal{A}}},$
@@
The term $f_{\mathcal{O}}(\hat{X})$ represents the goal of fitting the observation information, which evolves as the observation set is updated. The term $g(\hat{X})$ represents a kind of normalization of estimated states, aiming to conform the system structure.
Remark 2.
If the function $h(x)$ is an $L$-smooth and $\lambda$-strongly convex function, it satisfies the following condition: $\forall x,x^{0}$,
@@
Suppose the smooth and strong convexity of $f_{\mathcal{O}}(\hat{X})$ and $g(\hat{X})$ holds with $L_{f},\lambda_{f},L_{g},\lambda_{g}$, then the following inequalities hold
$L_{f}$ =$=\begin{cases}0,&\text{if }\mathcal{O}=\emptyset\\ \lambda_{\text{max}}(C^{T}R^{-1}C),&\text{otherwise}\end{cases},$
$\lambda_{f}$ =$=\begin{cases}\lambda_{\text{min}}(C^{T}R^{-1}C),&\text{if }\mathcal{O}=\{0,1,2,\dots,N\}\\ 0,&\text{otherwise}\end{cases},$
@@
Therefore, we can use proximal gradient descent to find the optimal solution iteratively. Define the function $f^{t}_{\mathcal{O}}(x)$ as the approximate function, and $prox_{\eta g}(\cdot)$ as the proximal operator:
$f^{t}_{\mathcal{O}}(x)=f(x^{t})+\nabla f_{\mathcal{O}}(x^{t})^{T}(x-x^{t})+\frac{1}{2\eta_{t}}\|x-x^{t}\|^{2}_{2},$ (16)
$prox_{\eta g}(x)=\arg\min_{z}\left[\frac{1}{2}\|z-x\|^{2}_{2}+\eta g(z)\right].$ (17)
Theorem 2 demonstrates how to iteratively solve for the optimal solution to Problem 4. The proof can be found in Appendix A-B.
@@
Suppose $X^{*}$ is the unique solution of Problem 4. With the fixed learning rate $\eta=\frac{1}{\lambda_{\text{max}}(C^{T}R^{-1}C)}\leq\frac{1}{L_{f}}$, repeat the following step and $X^{t}$ will converge to $X^{*}$.
@@
We can interpret this iterative process from the perspective of Kalman filter and RTS smoother. The term $X^{t}-\eta\nabla f_{\mathcal{O}}(X^{t})$ is comparable to the measurement update step in Kalman filtering, where $\nabla f_{\mathcal{O}}(X^{t})$ corresponds to the residual and a fixed parameter $\eta$ corresponds to the Kalman gain. The term $(I+\eta H)^{-1}$ plays the role of a smoother, updating the estimate based on the system structure.
Solving Problem 4 using Theorem 2 will yield the same result as solving Problem 3 with Lemma 1. However, compared to Lemma 1, the iterative approach provides by Theorem 2 does not require calculating the Kalman gain and smoother gain at each step. Instead, it replaces the Kalman gain with a fixed parameter $\eta$ and the smoother gain with a fixed matrix $(I+\eta H)^{-1}$, converging to the optimal estimate after several iterations. This provides a new perspective on estimation update.
V-B Estimation update for the disordered new observation
@@
Input: the system parameters $\{A,C,Q,R\}$ and the stop condition $\epsilon$
Input: initial state $(\bar{x}_{0}=0,\bar{P}_{0})$ and the full observation set $y_{0:N}$
Output: estimation of state sequence $\{\hat{x}_{0},\cdots,\hat{x}_{N}\}$
1 Calculate the matrix $H,L_{f}$ and $R^{-1}$;
2 Choose the $\eta<1/L_{f}$ and calculate $(I+\eta H)^{-1}$;
3 Initialize $\mathcal{O}=\emptyset$, $\hat{X}_{\mathcal{O}}=col\{\hat{x}_{0|\mathcal{O}}=0,\cdots,\hat{x}_{N|\mathcal{O}}=0\}$;
// update with new observations
4 repeat
5       if $\mathcal{O}$ changes to $\mathcal{O}^{*}$ then
6             $\hat{X}=\hat{X}_{\mathcal{O}}$;
7             repeat
8                   $\nabla f_{\mathcal{O}^{*}}=\sum_{i\in\mathcal{O}^{*}}C_{i}^{T}R^{-1}(y_{i}-C_{i}\hat{X})$;
9                   $\hat{X}^{*}=(I+\eta H)^{-1}(\hat{X}-\eta\nabla f_{\mathcal{O}^{*}}(\hat{X}))$;
10                   if $|\mathcal{L}_{\mathcal{O}^{*}}(\hat{X})-\mathcal{L}_{\mathcal{O}^{*}}(\hat{X}^{*})|<\epsilon$ then
11                         break;
12
13                  else
14                         $\hat{X}=\hat{X}^{*}$;
15
16
17            until ;
18            $\mathcal{O}=\mathcal{O}^{*},\hat{X}_{\mathcal{O}}=\hat{X}^{*}$;
19
20
21until ;
@@ 18            $\mathcal{O}=\mathcal{O}^{*},\hat{X}_{\mathcal{O}}=\hat{X}^{*}$;
Compared to the direct approach, this iterative method for updating estimates has two advantages. Firstly, the iterative method does not compute the inverse matrix to obtain the optimal gain, but only requires matrix multiplication, and the fixed inverse matrices $(I+\eta H)^{-1}$ and $R^{-1}$ can be calculated in advance. Secondly, the previous estimate can be used as the initial value of the iteration, and the information in the previous observation set can be fully utilized.
When $|y_{\mathcal{O}}|$ is large, the term $f_{\mathcal{O}}$ has slightly change with new observations, leading to minor differences between $\hat{X}_{\mathcal{O}^{*}}$ and $\hat{X}_{\mathcal{O}}$. This implies that, as the impact of the new observation on $f_{\mathcal{O}}$ becomes more negligible, the algorithm will converge more rapidly. Theorem 3 provides a quantitative analysis of the maximum number of iterations when new observations are added. The proof can be found in Appendix A-C.
@@
$\tau<\log_{1-\theta}\left({\frac{\epsilon}{\|\Xi\xi\|^{2}_{R^{-1}}+2\|\Xi\|_{R^{-1}}\|\xi\|_{R^{-1}}}}\right)$ (20)
@@
Theorem 3 demonstrates that the more reliable the initial $\hat{x}_{k}$ is, the fewer iterations are required for convergence. Thus, as the observation set $y_{\mathcal{O}}$ expands, the number of iterations required for the estimation update decreases.
VI Algorithm and analysis
In this section, we aim to solve the targeted MIP problem presented in (7). Summarized in Fig. 6, the GBS estimator involves repeating two steps: assessing the reliability of observations and updating estimates. The algorithms introduced in Section IV and Section V facilitate the estimation update in response to changes in reliable observation set.
@@ VI Algorithm and analysis
VI-A Initial strategy for rough estimation
A good initial value is beneficial in solving problems. However, at the beginning, we do not know the initial values of $\hat{p}_{0:N}$ and $\hat{x}_{0:N}$. Therefore, the initial strategy is targeted to roughly estimate the value of observation indicators $p_{0:N}$ based on $y_{0:N}$.
@@ VI-A Initial strategy for rough estimation
Input: the system parameters $\{A,C,Q,R,\alpha\}$
Input: the initial state estimate $(\bar{x}_{0}=0,\bar{P}_{0})$ and the full observation set $y_{0:N}$
Output: estimated indicators $\hat{p}_{0:N}$ and state $\hat{x}_{0:N}$
1 Initial $\hat{x}_{0}=\bar{x}_{0},P_{0}=\bar{P}_{0}$;
2 for $i=0$ to $N$ do
       // determine the indicators
3       $\hat{p}_{i}=\textbf{bool}(\|y_{i}-C\hat{x}_{i}\|^{2}_{R^{-1}}<\alpha)$;
       // estimate the state
4       if $p_{i}=1$ then
5             $K_{i}=P_{i}C^{T}(CP_{i}C^{T}+R)^{-1}$;
6             $\hat{x}_{i}=\hat{x}_{i}+K_{i}(y_{i}-C\hat{x}_{i})$;
7             $P_{i}=(I-K_{i}C)P_{i}$;
8
9       $\hat{x}_{i+1}=A\hat{x}_{i}$;
10       $P_{i+1}=AP_{i}A^{T}+Q$;
11
@@ 10       $P_{i+1}=AP_{i}A^{T}+Q$;
VI-B Gaussian-Bernoulli Secure (GBS) estimator
The approach to solving Problem 1 is to iteratively seek the optimal estimate by fixing the $\hat{p}_{0:N}$ and $\hat{x}_{0:N}$, respectively.

### v2

@@ 4.2 Estimation Update with New Observations
A basic approach is illustrated in Fig. 5 by Lemma 1. In detail, we utilize the Kalman gain $\{K_{t}|t<k\}$ under $y_{\mathcal{O}}$ and recompute the rest process under $y_{\mathcal{O}^{*}}$.
Figure 5: Recalculation steps for estimation update.
However, employing a direct method to compute the optimal estimate reveals limitations in computational cost. When $|y_{\mathcal{O}}|$ is large, especially in systems with multiple sensors, the direct method requires computing the matrix inverse $O(|y_{\mathcal{O}}|)$ times, which consumes significant computational resources. Furthermore, as the observation set expands, the optimal state estimate approaches the real state more closely. Nevertheless, adding new observations to a larger set $y_{\mathcal{O}}$ requires an increasing number of calculation steps to update the estimate, especially in multi-sensor systems. This indicates that using direct method to update cannot fully leverage the information from the previous estimation, and thus causes a waste of computational resources.
5 State Estimation with Secure Data: An Iterative Computational Approach
In section 4, we discussed how to find the closed-form solution of equation (8), but its computational efficiency is poor. In this section, we propose an iterative method for state sequence estimation that allows for rapid convergence to the optimal estimate.
The core idea of the iterative method is to fully utilize the previous estimates as the initial guess in the iteration. By constructing a composite convex optimization problem, we design a proximal gradient descent algorithm to avoid redundant computations. Furthermore, we derive the convergence rate and the iteration count through the reliability of previous estimates.
5.1 Sequential Estimation with Partial Observations
The objective function in equation (8) incorporates two goals: first, the state estimates $\hat{x}_{0:N}$ should match the observations $\hat{y}_{\mathcal{O}}$; second, the state estimates $\hat{x}_{0:N}$ should conform to the known structure of the system. Therefore, $\mathcal{L}_{\mathcal{O}}$ can be considered as the sum of two functions:
@@ 5.1 Sequential Estimation with Partial Observations
By stacking all $\hat{x}_{i}$ into a vector $\hat{X}=col\{\hat{x}_{0|\mathcal{O}},\cdots,\hat{x}_{N|\mathcal{O}}\}$, we construct a sparse matrix $\tilde{\mathcal{A}}$ to capture the relationship between the states:
$\begin{bmatrix}\hat{x}_{0}\\ \hat{x}_{1}-A\hat{x}_{0}\\ \hat{x}_{2}-A\hat{x}_{1}\\ \vdots\\ \hat{x}_{N}-A\hat{x}_{N-1}\end{bmatrix}=\begin{bmatrix}I&0&0&\cdots&0\\ -A&I&0&\cdots&0\\ 0&-A&I&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&I\end{bmatrix}\begin{bmatrix}\hat{x}_{0}\\ \hat{x}_{1}\\ \hat{x}_{2}\\ \vdots\\ \hat{x}_{N}\end{bmatrix}=\tilde{\mathcal{A}}\hat{X}$.
@@
$\tilde{P}=\begin{bmatrix}~~P_{0}~~&0&0&\cdots&0\\ 0&~~Q~~&0&\cdots&0\\ 0&0&~~Q~~&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&~~Q~~\end{bmatrix}$.
With the help of equation (12), we convert Problem 2 to following composite convex optimization problem.
@@
[…]
@@
[…]
@@
[…]
Definition 1.
The function $h(x)$ is an $L$-smooth and $\lambda$-strongly convex function if it satisfies the following condition: $\forall x,x^{0}$,
@@
[…]
$L_{f}$ $=\begin{cases}0,&\text{if }\mathcal{O}=\emptyset\\ \lambda_{\text{max}}(C^{T}R^{-1}C),&\text{otherwise}\end{cases},$
$\lambda_{f}$ $=\begin{cases}\lambda_{\text{min}}(C^{T}R^{-1}C),&\text{if }\mathcal{O}=\{0,1,2,\dots,N\}\\ 0,&\text{otherwise}\end{cases},$
@@
[…]
In addition, we use proximal gradient descent to iteratively optimize the solution. Define $f^{t}_{\mathcal{O}}(x)$ as the approximate function and $prox_{\eta g}(\cdot)$ as the proximal operator:
$f^{t}_{\mathcal{O}}(x)$ $=f(x^{t})+\nabla f_{\mathcal{O}}(x^{t})^{T}(x-x^{t})+\frac{1}{2\eta_{t}}\|x-x^{t}\|^{2}_{2},$
$prox_{\eta g}(x)$ $=\arg\min_{z}\left[\frac{1}{2}\|z-x\|^{2}_{2}+\eta g(z)\right].$
The following theorem demonstrates how to iteratively solve for the optimal solution to Problem 4.
@@
Suppose $X^{*}$ is the unique solution of Problem 4. With the fixed learning rate $\eta=\frac{1}{\lambda_{\text{max}}(C^{T}R^{-1}C)}\leq\frac{1}{L_{f}}$, repeat the following step and $X^{t}$ will converge to $X^{*}$:
@@
We can interpret this iterative process from the perspective of filter and smoother. The term $X^{t}-\eta\nabla f_{\mathcal{O}}(X^{t})$ is comparable to the measurement update step in Kalman filtering, where $\nabla f_{\mathcal{O}}(X^{t})$ corresponds to the residual and a fixed parameter $\eta$ corresponds to the Kalman gain. The term $(I+\eta H)^{-1}$ can be regarded as a smoother, updating the estimate based on system dynamics.
Solving Problem 4 using Theorem 2 will yield the same result as solving Problem 3 with Lemma 1. However, compared to Lemma 1, the iterative approach provides by Theorem 2 does not require calculating the Kalman gain and smoother gain at each step. Instead, it replaces the filter gain with a fixed parameter $\eta$ and the smoother gain with a fixed matrix $(I+\eta H)^{-1}$, converging to the optimal estimate after several iterations. This provides a new perspective on estimation update.
5.2 Estimation Update with New Observations
@@
1:system parameters $\{A,C,Q,R\}$, stop condition $\epsilon$
2:initial state $(\bar{x}_{0}=0,\bar{P}_{0})$, full observation set $y_{0:N}$
3:estimation of state sequence $\{\hat{x}_{0},\cdots,\hat{x}_{N}\}$
4:Calculate the matrix $H$, Lipschitz constant $L_{f}$, and $R^{-1}$
5:Choose $\eta<1/L_{f}$ and compute $(I+\eta H)^{-1}$
6:Initialize $\mathcal{O}\leftarrow\emptyset$, $\hat{X}_{\mathcal{O}}\leftarrow\text{col}\{\hat{x}_{0|\mathcal{O}}=0,\cdots,\hat{x}_{N|\mathcal{O}}=0\}$
[…]
8:  if $\mathcal{O}$ changes to $\mathcal{O}^{*}$ then
9:    $\hat{X}\leftarrow\hat{X}_{\mathcal{O}}$
10:    repeat
[…]
12:     $\hat{X}^{*}\leftarrow(I+\eta H)^{-1}(\hat{X}-\eta\nabla f_{\mathcal{O}^{*}})$
13:     if $|\mathcal{L}_{\mathcal{O}^{*}}(\hat{X})-\mathcal{L}_{\mathcal{O}^{*}}(\hat{X}^{*})|<\epsilon$ then
14:      break
15:     else
16:      $\hat{X}\leftarrow\hat{X}^{*}$
17:     end if
18:    until converged
19:    $\mathcal{O}\leftarrow\mathcal{O}^{*}$, $\hat{X}_{\mathcal{O}}\leftarrow\hat{X}^{*}$
20:  end if
21:until $\mathcal{O}$ remains unchanged
Compared to the direct approach, this iterative method for updating estimates has two advantages. Firstly, the iterative method does not compute the inverse matrix to obtain the optimal gain, but only requires matrix multiplication, and the fixed inverse matrices $(I+\eta H)^{-1}$ and $R^{-1}$ can be calculated in advance. Secondly, the previous estimate can be used as the initial value of the iteration and the information from the previous observation set can be fully utilized.
When $|y_{\mathcal{O}}|$ is large, the term $f_{\mathcal{O}}$ changes slightly with new observations, leading to minor differences between $\hat{X}_{\mathcal{O}^{*}}$ and $\hat{X}_{\mathcal{O}}$. This implies that as the impact of the new observation on $f_{\mathcal{O}}$ becomes more negligible, the algorithm will converge more rapidly. Theorem 2 provides a quantitative analysis of the maximum number of iterations when new observations are added.
@@
$\tau<\log_{1-\theta}\left({\frac{\epsilon}{\|\Xi\xi\|^{2}_{R^{-1}}+2\|\Xi\|_{R^{-1}}\|\xi\|_{R^{-1}}}}\right),$ (17)
@@
[…]
Theorem 2 implies that the more reliable the initial $\hat{x}_{k}$ is, the fewer iterations are required for convergence. Thus, as the observation set $y_{\mathcal{O}}$ expands, the number of iterations required for the estimation update decreases.
6 Secure State Estimation Method and Convergence Analysis
In this section, we aim to solve the targeted MIP problem presented in (7). Summarized in Fig. 6, the GBS estimator involves repeating two steps: assessing the reliability of observations and updating estimates. The algorithms introduced in Section 4 and Section 5 facilitate the estimation update in response to changes in reliable observation set.
@@ 6 Secure State Estimation Method and Convergence Analysis
6.1 Initial Strategy for Rough Estimation
A good initial value is beneficial for solving problems. However, at first, we do not know the initial values of $\hat{p}_{0:N}$ and $\hat{x}_{0:N}$. Therefore, the initial strategy aims to roughly estimate the value of the observation indicators $p_{0:N}$ based on $y_{0:N}$.
@@ 6.1 Initial Strategy for Rough Estimation
1:system parameters $\{A,C,Q,R,\alpha\}$
2:initial state estimate $(\bar{x}_{0}=0,\bar{P}_{0})$ and full observation set $y_{0:N}$
3:estimated indicators $\hat{p}_{0:N}$ and states $\hat{x}_{0:N}$
4:Initialize $\hat{x}_{0}\leftarrow\bar{x}_{0}$, $P_{0}\leftarrow\bar{P}_{0}$
5:for $i=0$ to NN do
6:  $\hat{p}_{i}\leftarrow\textbf{bool}(\|y_{i}-C\hat{x}_{i}\|^{2}_{R^{-1}}<\alpha)$
7:  if $\hat{p}_{i}=1$ then
8:    $K_{i}\leftarrow P_{i}C^{T}(CP_{i}C^{T}+R)^{-1}$
9:    $\hat{x}_{i}\leftarrow\hat{x}_{i}+K_{i}(y_{i}-C\hat{x}_{i})$
10:    $P_{i}\leftarrow(I-K_{i}C)P_{i}$
11:  end if
12:  $\hat{x}_{i+1}\leftarrow A\hat{x}_{i}$
13:  $P_{i+1}\leftarrow AP_{i}A^{T}+Q$
14:end for
6.2 Gaussian-Bernoulli Secure (GBS) Estimator
The approach to solving Problem 1 is to iteratively optimize the estimates by fixing $\hat{p}_{0:N}$ and $\hat{x}_{0:N}$, respectively.